woody woodpecker pecked at a 17 meter wooden pole until it cracked and the upper part fell with the top hitting the ground 10m from the foot of the pole. since the upper part had not completely broken off, woody pecked away where the pole had cracked. how far was woody above the ground?
Answer by reelmccray(6)

For triangles
A = Height
B = Base
C = Hypotenuse
A^2 + B^2 = C^2
In this case B = 10
A + C = 17 meters
Which can be written as:
A = -C + 17
Substitute (-C + 17) in A^2 + B^2 = C^2
(-C + 17)^2 + 10^2 = C^2
Use FOIL to figure out (-C + 17)^2
(C^2 - 34C + 289) + 100 = C^2
C^2 - 34C + 389 = C^2
34C = 389
C = 11.44
A + C = 17 so A + 11.44 = 17
A = 5.56 meter
Woody is 5.56 meters off the ground